DETERMINATION OF THE 0_{g}^{−} PURE LONG RANGE POTENTIAL CURVE OF Rb_{2} AND Cs_{2}; APPLICA-TION TO ULTRACOLD MOLECULES FORMATION

\underline{R. F. GUTTERRES}, C. AMIOT, O. DULIEU and F. MASNOU-SEEUWS, Laboratoire Aimé Cotton, Bât 505, Campus d’Orsay 91405, Orsay Cedex, France.

This work presents an accurate study of the 0_{g}^{−}(P_{3/2}) pure long range electronic state of the ^{87}Rb(5S)-^{87}Rb(5P_{3/2}) and 133Cs(6S)-^{133}Cs(6P_{3/2}) molecular systems. Generalized simulated annealing method was used in order to reduce the high resolution spectral data provided by photoassociative spectroscopy of ultra cold ^{87}Rb and ^{133}Cs atoms. The analysis has allowed the determination of the effective dispersion parameters, the chemical exchange energy contribution as well as the dissociation energy concerning the analytical potential representation of the 0_{g}^{−}(P_{3/2}) state for both, Rb_{2} and Cs_{2}, molecules. A detailed comparison with the potential curves obtained through RKR procedure is performed, and the consequences for ultracold molecules formation rates are discussed.